# Extending 1089 attractor to any number of digits and any number of steps

### v1

@@ ABSTRACT
The well-known “1089” trick reflects an amazing trait of digital reversal process and reminisces of a limiting attractor in dynamical systems even though it takes only two steps. It is natural to consider the situations when the number of digits is beyond three as in the original 1089 trick, as well as situations when the number of steps is beyond two. The generalization to a larger number of digits has been mostly done by Webster’s work which we will reproduce. After the two steps of the “1089” trick for any number of digits, the resulting integers, the number of which is very low, are named “Papadakis-Webster integers” (PWI). A PWI is always divisible by 99, and the resulting quotients consist of only 0’s and 1’s, which we name “Papadakis-Webster binary strings” (PWBS). Not all binary strings could be PWBS, and we define the “hairpin pairing rule” to determine if a binary string is a PWBS. To generalize 1089 trick to any number of steps, we propose a two-option iteration system named “iterative digital reversal” (IDR) suitably interweaving additions and subtractions. The simplest limiting behavior of IDR is 2-cycles. The elements in an IDR 2-cycle are all composed of repetitions of the 10(9)L89 ($L\geq 0$) motif, and are all PWIs. The lower 2-cycle elements after division of 99 belong to the subset of PWBS that are palindromic and consist of 0- and 1-blocks with a minimal length of two. IDR also has higher $p$-cycles ($p=$ 10, 12, 71) whose elements seem to contain at least one PWI. Another interesting finding about IDR is that it contains non-periodic and diverging trajectories, as the integer values grow to infinity. In these diverging trajectories, while the number of flanking digits around the middle point increases by the iteration, the middle part has an 8-cycle rhythm or signature which has been found in all diverging trajectories. Overall, the generalization of the original 1089 trick in both “space” and “time” leads to many new patterns in integers and new phenomenology in dynamics.
@@ 1 Introduction
In an even earlier work, Constantinos Papadakis, a Greek engineer and inventor, also attempted to generalize the 1089 trick to any number of digits (Papadakis, 1982). He described the properties of what we call here Papadakis-Webster numbers and binary strings and emphasized the significance of their very low populations. Still, his proofs were somewhat amateurish and incomplete. His work is even less known because his monograph is not in English (for historical reasons, we plan to make his monograph freely available on the Internet soon, with an introduction in English).
@@
To our surprise, 2-cycle attractors are not the only possible end results of our operation with infinite number of steps. Other longer cycles are possible. To our surprise, there are even non-periodic (acyclic) behaviors: the integers become larger and larger as we continue to iterate the mapping, even though subtraction part of the operation is always used. There are constraints on the integers in the diverging trajectories. There are also 8-cycle rhythm or “signature” in the digital sequence of these non-attractor integers. All these results are well beyond the original 1089 trick and Webster’s work.
The paper is organized as follows: section 2 is on extending 1089 trick to larger integers with more than 3 digits, principally studying the particular properties of the (very few) end-integers of the procedure. It has two subsections. One is about the produced end-integers themselves, the Papadakis-Webster integers. The other is about the Papadakis-Webster integers (PWI) divided by 99, which always end up to integers that consist of 0’s and 1’s only, called Papadakis-Webster binary strings (PWBS). Section 3 is about extending 1089 trick to infinite number of steps which contains five subsections. The first subsection introduces the iterative digital reversal (IDR) mapping and the general description of its dynamical behavior. The second and the third subsections describe the 2-cycle attractors of IDR and its connection to the Papadakis-Webster integers. The fourth and fifth subsections are about higher cyclic attractors, and acyclic diverging trajectories. The paper ends with the Discussion section. The Appendices contain discussion on a caveat in the 1089 trick, on the requirement that the number of digits after subtraction has to remain the same; more examples of PWI not listed in the main text; and an example of a 71-cycle.
@@ 2.1 Introducing the Papadakis-Webster integers and Papadakis-Webster binary strings
    (Digital reversal): For any integers with $n+1$ digits, $D=\sum_{i=0}^{n}a_{i}10^{i}=(a_{n}a_{n-1}\cdots a_{2}a_{1}a_{0})$, its digital reverse $rev(D)$ is defined as $rev(D)\equiv\sum_{i=0}^{n}a_{n-i}10^{i}=(a_{0}a_{1}a_{2}\cdots a_{n-1}a_{n})$.
@@
Table 1: All PWI and the corresponding PWBS when the starting integer’s length (n+1, for $a_{n}a_{n-1}\cdots a_{1}a_{0}$) is 1-7. Binary strings that are not PWBS are also listed (last column). The number of PWI’s as a function of $n$ follows a (partial) Fibonacci sequence (1,1,3,3,8,8,21,21…). The PWI and the corresponding PWBS for n+1=8,9 are included in the Appendix.
@@ 2.2 Hairpin pairing rule for Papadakis-Webster binary strings
Figure 1: Illustration of the hairpin pairing rule for PWBS’s. To check if a binary string is PWBS, remove the leading 1 from the sequence, then fold the rest of the sequence around its middle position (either between two middle digits, or the middle digits itself). If all pairing digits have different values, it is a PWBS. If pairing digits are both 1’s, as long as their left digits are also both 1’s, it is fine; similarly, if pairing digits are both 0’s, it will be find if the digits on their left are also both 0’s. For the self-pairing digit in the middle position, its left digit should be the same as itself in order to be PWBS. Using this rule, (B)(C)(D) are PWBS, whereas (A) is not PWBS. The double-head arrows mark the digits in pairing position that have the same value. The single-head arrows mark the self-pairing digits.
@@ 3.2 Constraints on the integers in the 2-cycle attractor
For two elements in the 2-cycle attractor of IDR, $D_{I+1}>D_{I}$, then $rev(D_{I})/D_{I}=9$.
Since $D_{I+1}=D_{I}+rev(D_{I})=D_{I}\times 10$ according to Proposition 3.1, $rev(D_{I})=9\times D_{I}$. The fact that 9801 is divisible by its digital reversal was mentioned in (Hardy, 1940).
@@
Outline of a proof of Proposition 3.6: similar to the proof of Proposition 3.4 where we show that the property of $D_{I+1}=10D_{I}$ forces the first two digits to be 0,1, and the last digits to be 8,9, we can continue to examine the constrain towards the middle of the sequence. For example, one can show that the first three and the last three digits can either (1,0,9$\cdots$,9,8,9), or (1,0,8$\cdots$,0,8,9). The first 4 and last 4 digits in the first situation would be (1,0,9,9, $\cdots$, 9,9,8,9), and in the second situation (1,0,8,9, $\cdots$, 1,0,8,9), etc. Once the uniqueness of $M_{L}=10(9)_{L}89$ as the fundamental 2-cycle motif is established, its symmetric concatenation with padding zeros in various forms can be shown also to be 2-cycles, similar to the proof for Proposition 3.5. In the Discussion section, we will mention that alternative proofs can be found in the literature (Hoey, 1992; Webster and Williams, 2012).
@@
Theorem 3.8
The binary strings described in Conjecture 3.7, i.e., symmetric arrangement of 0-block and 1-block, whose lengths are larger or equal to 2, are PWBS.
  • Proof
Theorem 3.8 shows that all limiting 2-cycle integers of IDR are PWI.
@@ 3.4 Construction of an infinite number of $p$-cycles ($p>2$)
Concatenations of $p$-cycle elements by padding certain number of zeros in-between, are also $p$-cycle elements. For example, Table 3(B) shows that combining two 12-cycle elements 99099s (as seen in Table 3(A)) with four zeros between them, 99099 $0$ 99099, is a 12-cycle member. Comparing integers in Table 3(A) and (B), each row can find its correspondence in another table, though the number of zero-padding might be different (e.g., 399049893 in Table 3(A) vs. 399049893 399049893 in Table 3(B)).
@@ 3.5 Non-periodic plus diverging trajectories with infinite number of steps
For a diverging/non-periodic trajectory of IDR, the middle digit(s) follow a 8-cycle rhythm: 98, 08, 8, 9, 99, 99, 9, and 9.
@@ 4 Discussion
While the extension of 1089 trick to any number of digits is straightforward, the extension to any number of steps need more discussion. Our Eq.9 aims at applying both subtraction and addition of an integer $D$ and its digital reverse $rev(D)$. However, Eq.9 is not unique. For example, suppose we change the condition for applying subtraction from $rev(D_{i})<D_{i}$ to $rev(D_{i})\leq D_{i}$:
@@
Adding or subtracting an integer’s digital reverse from the integer itself, especially if done in multiple or repetitive ways, is an operation leading to integers obeying to very specific structural restrictions. These include the digital composition, the appearance of specific digital motifs, as well as the divisibility properties of the outcome. Therefore, the resulting integers, either after only two steps or after infinite number of steps, could fall into a very small set in the $\mathbb{N^{+}}$ space. Indeed, the proportion of integers that are PWI decreases exponentially with the number of digits. The limiting 2-cycle elements of IDR are a subset of PWI, whose number is even fewer. Even though there is an infinite number of elements in a diverging trajectory, percentage-wise these still occupy a very small set since there are specific signature pattern in these integer sequences.
@@
Corollary 3.3 states that if $D_{I}$ is the smaller element in a limiting 2-cycle set of IDR, then $rev(D_{I})/D_{I}=9$. Integers with this property (digital reversal of an integer being divisible by that integer) is called “reverse divisible number” or “palintiple”, and has been investigated in multiple generations (Hardy, 1940; Beech, 1990; Webster and Williams, 2012; Holt, 2014; Kendrick, 2015; Holt, 2016). As we learned in Subsection 3.3, the 2-cycle elements all have a form of symmetric arrangement of the basic 10(9)L89 motifs. These 9-palintiples were actually independently discovered in https://oeis.org/A008919 , and the exactly identical form of 9-palintiple as our Propositions 3.5 and 3.6 was presented in (Webster and Williams, 2012).
Although all the results in this article concern decimal integers (i.e., with base 10), similar results can be obtained for non-decimal bases. For example, as already pointed out by Webster (Webster, 1995), the divisibility by 99 for decimal PWI will become divisibility (B+1)(B-1) (where B is the base) for non-decimal PWIs. For example, if B=2, non-decimal PWIs are always multiples of 3, if B=8, they are multiples of 63, etc. New cycle lengths have also been observed for non-decimal IDR (results not shown).
Many mathematical programs or calculators can not carry out arithmetics correctly for very larger integers. We have observed, for example, that the R software environment (https://www.r-project.org/) could produce incorrect output for simple division of a very large number. To aid such large number arithmetics, we provide a FORTRAN source code IDR1f.f90, detailed guidelines, and its executables in Windows environment at:
https://shorturl.at/PkwJ6
https://shorturl.at/0Qje5
https://shorturl.at/G1g1y
@@
Acknowledgement: We would like to thank Dr. Astero Provata for helpful discussions and reading the manuscript. YA thanks Dr. Marianna Vasilakaki for computer assistance, and Ms. Maria Metaxa for sharing a copy of C. Papadakis’ monograph. YA would like to dedicate this work to the memory of Michael Rossos – a friend, an old gentleman, a polymath with a special interest in linguistics and ancient religions. It was him, after reading a mathematics popularization article about Papadakis’ work, who started to experiment with numbers adding or subtracting them after reversion, conditioning on their relative values. He was already amazed by how often number 1089 appeared in the output. Through his playful creative activity, he was the true inventor of IDR almost 30 years ago.

### v2

@@ ABSTRACT
The well-known “1089” trick reflects an amazing trait of digital reversal process and reminisces of a limiting attractor in dynamical systems even though it takes only two steps. It is natural to consider the situations when the number of digits is beyond three as in the original 1089 trick, as well as situations when the number of steps is beyond two. The generalization to a larger number of digits has been mostly done by Webster’s work which we will reproduce. After the two steps of the “1089” trick for any number of digits, the resulting integers, the number of which is very low, are named here “Papadakis-Webster integers” (PWI). A PWI is always divisible by 99, and the resulting quotients consist of only 0’s and 1’s, which we name “Papadakis-Webster binary strings” (PWBS). Not all binary strings could be PWBS, and we define the “hairpin pairing rule” to determine if a binary string is a PWBS. To generalize 1089 trick to any number of steps, we propose a two-option iteration procedure named “iterative digital reversal” (IDR) suitably interweaving additions and subtractions. The simplest limiting behavior of IDR is 2-cycles. The elements in an IDR 2-cycle are all composed of repetitions of the 10(9)L89 ($L\geq 0$) motif, and are all PWIs. The lower 2-cycle elements after division of 99 belong to the subset of PWBS that are palindromic and consist of 0- and 1-blocks with a minimal length of two. IDR also has longer $p$-cycles ($p=$ 10, 12, 71) whose elements seem to contain at least one PWI. Another interesting finding about IDR is that it contains non-periodic and diverging trajectories, as the integer values grow to infinity. In these diverging trajectories, while the number of flanking digits around the middle point increases by the iteration, the middle part has an 8-cycle rhythm or signature which has been found in all diverging trajectories. Overall, the generalization of the original 1089 trick in both “space” and “time” leads to many new patterns in integers and new phenomenology in dynamics.
@@ 1 Introduction
In an even earlier work, Constantinos Papadakis, a Greek engineer and inventor, also attempted to generalize the 1089 trick to any number of digits (Papadakis, 1982). He described the properties of what we call here Papadakis-Webster numbers and binary strings and emphasized the significance of their very low populations. Still, his proofs were somewhat amateurish and incomplete. His work is even less known because his monograph is not in English. For historical reasons, we made his monograph freely available on the Internet, with an introduction in English, at https://shorturl.at/ILGdC.
@@
To our surprise, 2-cycle attractors are not the only possible end results of our operation with infinite number of steps. Other longer cycles are possible. Also, there are even non-periodic (acyclic) behaviors: the integers become larger and larger as we continue to iterate the mapping, even though subtraction part of the operation is always used. There are constraints on the integers in the diverging trajectories. There are also 8-cycle rhythm or “signature” in the digital sequence of these non-attractor integers. All these results are well beyond the original 1089 trick and Webster’s work.
The paper is organized as follows: section 2 is on extending 1089 trick to larger integers with more than 3 digits, principally studying the particular properties of the (very few) end-integers of the procedure. It has two subsections. One is about the produced end-integers themselves, the Papadakis-Webster integers. The other is about the Papadakis-Webster integers (PWI) divided by 99, which always end up to integers that consist of 0’s and 1’s only, called Papadakis-Webster binary strings (PWBS). Section 3 is about extending 1089 trick to an infinite number of steps which contains five subsections. The first subsection introduces the iterative digital reversal (IDR) mapping and the general description of its dynamical behavior. The second and the third subsections describe the 2-cycle attractors of IDR and its connection to the Papadakis-Webster integers. The fourth and fifth subsections are about higher cyclic attractors, and acyclic diverging trajectories. The paper ends with the Discussion section. The Appendices contain discussion on a caveat in the 1089 trick, on the requirement that the number of digits after subtraction has to remain the same; more examples of PWI not listed in the main text; an example of a 71-cycle; and a short discussion on the 1089 trick and IDR beyond decimal numerical system.
@@ 2.1 Introducing the Papadakis-Webster integers and Papadakis-Webster binary strings
    (Digital reversal): For any integers with $n+1$ digits, $D=\sum_{i=0}^{n}a_{i}10^{i}=(a_{n}a_{n-1}\cdots a_{2}a_{1}a_{0})$, its digital reversal $rev(D)$ is defined as $rev(D)\equiv\sum_{i=0}^{n}a_{n-i}10^{i}=(a_{0}a_{1}a_{2}\cdots a_{n-1}a_{n})$.
@@
Table 1: All PWI and the corresponding PWBS when the starting integer’s length (n+1, for $a_{n}a_{n-1}\cdots a_{1}a_{0}$) is 1-7. Binary strings that are not PWBS are also listed (last column). The number of PWI’s as a function of $n$ follows a (partial) Fibonacci sequence (1,1,3,3,8,8,21,21…). The PWI and the corresponding PWBS for n+1=8,9 are included in the Appendix A.2.
@@ 2.2 Hairpin pairing rule for Papadakis-Webster binary strings
Figure 1: Illustration of the hairpin pairing rule for PWBS’s. To check if a binary string is PWBS, remove the leading 1 from the sequence, then fold the rest of the sequence around its middle position (either between two middle digits, or the middle digit itself). If all pairing digits have different values, it is a PWBS. If pairing digits are both 1’s, as long as their left digits are also both 1’s, it is fine; similarly, if pairing digits are both 0’s, it will be fine if the digits on their left are also both 0’s. For the self-pairing digit in the middle position, its left digit should be the same as itself in order to be PWBS. Using this rule, (B)(C)(D) are PWBS, whereas (A) is not PWBS. The double-head arrows mark the digits in pairing position that have the same value. The single-head arrows mark the self-pairing digits.
@@ 3.2 Constraints on the integers in the 2-cycle attractor
For two elements in the 2-cycle attractor of IDR, $D_{I+1}>D_{I}$, then $rev(D_{I})/D_{I}=9$. Inversely, if rev(D)= 9D, then D and 10D are the two 2-cycle elements of IDR.
Since $D_{I+1}=D_{I}+rev(D_{I})=D_{I}\times 10$ according to Proposition 3.1, $rev(D_{I})=9D_{I}$. Similarly, if rev(D) = 9D, then D+rev(D) = 10D, 10 D-rev(D) = D, so D and 10 D are the 2-cycle elements of IDR. The fact that the digital reversal of 1089 is divisible by itself was mentioned in (Hardy, 1940).
  • Definition
    (Palintiple) If an integer D, whose digital reversal rev(D) is divisable by itself, i.e., rev(D)/D = k, where $k$ is a positive integer, then D is called a ($k$-)palintiple.
Corollary 3.3 is equivalent to the statements that (1) (the smaller member of) any 2-cycle of IDR is 9-palintiple; (2) any 9-palintiple integer, as well as its 10-multiple, are 2-cycle elements of IDR.
Note that other publications may define 9801 as palintiple. In our definition above, 1089 is a palintiple. Our definition is more convenient in the context of IDR.
@@
Outline of a proof of Proposition 3.6: similar to the proof of Proposition 3.4 where we show that the property of $D_{I+1}=10D_{I}$ forces the first two digits to be 0,1, and the last digits to be 8,9, we can continue to examine the constrain towards the middle of the sequence. For example, one can show that the first three and the last three digits can either (1,0,9$\cdots$,9,8,9), or (1,0,8$\cdots$,0,8,9). The first 4 and last 4 digits in the first situation would be (1,0,9,9, $\cdots$, 9,9,8,9), and in the second situation (1,0,8,9, $\cdots$, 1,0,8,9), etc. Once the uniqueness of $M_{L}=10(9)_{L}89$ as the fundamental 2-cycle motif is established, its symmetric concatenation with padding zeros in various forms can be shown also to be 2-cycles, similar to the proof for Proposition 3.5.
Alternatively, according to the Corollary 3.3, we only need to prove that the patterns described in Proposition 3.5 are the only integers for 9-palintiples. This proof can be found in (Hoey, 1992) (by Dan Hoey) and in (Webster and Williams, 2012) (by Roger Webster and Gareth Williams).
@@
Theorem 3.8
The binary strings described in Proposition 3.7, i.e., symmetric arrangement of 0-block and 1-block, whose lengths are larger or equal to 2, are PWBS.
  • Proof
Theorem 3.8 shows that all limiting 2-cycle integers of IDR are PWI.
@@ 3.4 Construction of an infinite number of $p$-cycles ($p>2$)
Concatenations of $p$-cycle elements by padding certain number of zeros in-between, are also $p$-cycle elements. For example, Table 3(B) shows that combining two 12-cycle elements 99099s (as seen in Table 3(A)) with four zeros between them, 99099 $0$ 99099, is a 12-cycle element. Comparing integers in Table 3(A) and (B), each row can find its correspondence in another table, though the number of zero-padding might be different (e.g., 399049893 in Table 3(A) vs. 399049893 399049893 in Table 3(B)).
@@ 3.5 Non-periodic plus diverging trajectories with infinite number of steps
There are an infinite number of diverging/non-periodic trajectoryies of IDR where the middle digit(s) follow a 8-cycle rhythm: 98, 08, 8, 9, 99, 99, 9, and 9.
@@ 4 Discussion
While the extension of 1089 trick to any number of digits is straightforward, the extension to any number of steps need more discussion. Our Eq.9 aims at applying both subtraction and addition of an integer $D$ and its digital reversal $rev(D)$. However, Eq.9 is not unique. For example, suppose we change the condition for applying subtraction from $rev(D_{i})<D_{i}$ to $rev(D_{i})\leq D_{i}$:
@@
Adding or subtracting an integer’s digital reverse from the integer itself, especially if done in multiple or repetitive ways, is an operation leading to integers obeying to very specific structural restrictions. These include the digital composition, the appearance of specific digital motifs, as well as the divisibility properties of the outcome. Therefore, the resulting integers, either after only two steps or after a very large number of steps, could fall into a very small set in the $\mathbb{N^{+}}$ space. Indeed, the proportion of integers that are PWI decreases exponentially with the number of digits. The limiting 2-cycle elements of IDR are a subset of PWI, whose number is even fewer. Even though there is an infinite number of elements in a diverging trajectory, percentage-wise these still occupy a very small set since there are specific signature pattern in these integer sequences.
@@
Corollary 3.3 states that if $D_{I}$ is the smaller element in a limiting 2-cycle set of IDR, then $rev(D_{I})/D_{I}=9$, i.e., a 9-palintiple. This connection between palintiples (Hardy, 1940; Sutcliffe, 1966; Beech, 1990; Pudwell, 2007; Holt, 2014; Kendrick, 2015; Holt, 2016) and 2-cycle of IDR provides another path in discovering sequence pattern of the integers in the limiting set. In fact, the motif 10(9)L89 and its extensions revealed in (Hoey, 1992; Webster and Williams, 2012) are exactly what we observed in numerical experiment of IDR.
Although all the results in this article concern decimal integers (i.e., with base 10), similar results can be obtained for non-decimal bases. For example, as already pointed out by Webster (Webster, 1995), the divisibility by 99 for decimal PWI will become divisibility (b+1)(b-1) (where b is the base) for non-decimal PWIs. For example, if b=2, non-decimal PWIs are always multiples of 3, if b=8, they are multiples of 63, etc. New cycle lengths have also been observed for non-decimal IDR (results not shown). More results for 1089 trick and IDR on base-b numerical system are included in Appendix A.4.
Many mathematical programs or calculators can not carry out arithmetics correctly for very larger integers. We have observed, for example, that the R software environment (https://www.r-project.org/) could produce incorrect output for simple division of a very large number. To aid such large number arithmetics, we provide a FORTRAN source code IDR1f.f90, detailed guidelines, and its executable in Windows environment at: https://shorturl.at/PkwJ6 , https://shorturl.at/0Qje5 , https://shorturl.at/G1g1y .
@@
Acknowledgement: We would like to thank Dr. Astero Provata for helpful discussions and reading the manuscript. YA thanks Dr. Marianna Vasilakaki for computer assistance, and Ms. Maria Metaxa for sharing a copy of C. Papadakis’ monograph. WL thanks Oliver Clay for helpful comments on the draft. YA would like to dedicate this work to the memory of Michael Rossos – a friend, an old gentleman, a polymath with a special interest in linguistics and ancient religions. It was him, after reading a mathematics popularization article about Papadakis’ work, who started to experiment with numbers adding or subtracting them after reversion, conditioning on their relative values. He was already amazed by how often number 1089 appeared in the output. Through his playful creative activity, he was the true inventor of IDR almost 30 years ago.
@@ Appendix A.3 Integers in a limiting 71-cycle
Appendix A.4 IDR beyond the decimal system
A base-b length-n integer is defined as $D=\sum_{i=0}^{n}a_{i}b^{i}=(a_{0}a_{1}a_{2}\cdots a_{n})$, where $b>1$ is a positive integer, and $a_{i}\in(0,1,2,\cdots,b-2,b-1)$. The digital reverse of $D$ is defined as before: $rev(D)=(a_{n}a_{n-1}\cdots a_{2}a_{1}a_{0})$. The Papadakis-Webster integers (PWI) in base-b system are defined the same: E $\equiv$ D-rev(D), PWI $\equiv$ E+rev(E). The mapping of Eq.9 (IDR) is also defined the same as before. Here, we describe some results concerning 1089 trick and IDR, generalized from decimal system to any base-b systems, omitting proofs. We restrict ourselves to systems with $b>2$, as the binary system ($b=2$) presents certain specific peculiarities.
  • The concept of divisibility for base-b integers is based on nodulo arithmetic (Uspensky and Heaslet, 1939). Theorem 2.2 that PWI is divisible by 99 for decimal system is generalized to base-b system as: base-b PWI is divisible by (b-1)(b+1), and the quotient is a binary string. Therefore, the concept of Papadakis-Webster binary string (PWBS) remains to be true.
  • Proposition 3.1 and corollary 3.3 is generalized to: for base-b integers in a limiting 2-cycle of IDR, $D_{I+1}=bD_{I}$, and $rev(D_{I})/D_{I}=b-1$.
  • The propositions 3.5 and 3.6, concerning 10(9)L89 ($L\geq 0$) motifs and its expansions as 2-cycle elements for IDR, can be generalized to 10[b-1]L[b-2][b-1] ($L\geq 0$, where [b-1] and [b-2] are the symbols presenting values b-1 and b-2) motifs, as 2-cycle elements for base-b integers. For example, for base-8 integers, the foundamental motif is 10(7)L67; for base-13 integers (where the digits are 0,1,2, $\cdots$, 9, a,b,c), the motif is 10(c)Lbc.
  • For limiting $p$-cycle of IDR, same cycle length may also appear in other base-b systems with a similar pattern. See Table A.3, for example of a $p=12$ cycle in base-8 (octal) and base-9 (nonal) integers, as compared to the correponding decimal system.
  • There are also $p$-cycles in base-b integer systems that do not have a correspondence in the decimal system. Table A.4 shows such an example of a 41-cycle in base-4 (quaternary) system.
  • Similar to Conjecture 3.12 that diverging trajectories of IDR tend to have a 8-cycle rhythm in the middle section for the decimal system, there is also a ubiquitous 8-cycle rhythm in the middle digits in base-b numerical systems: [b-1][b-2], 0[b-2], [b-2], [b-1], [b-1][b-1], [b-1][b-1], [b-1],and [b-1].
An example of 12-cycle of IDR in base-8 and base-9 integers as compared to the decimal system
i b=8 b=9 b=10
1 77077 88088 99099
2 176176 187187 198198
3 1070067 1080078 1090089
4 10670770 10780880 10890990
5 761167 871178 981189
6 1742356 1852367 1962378
7 10475047 10585058 10695069
8 104554450 105654560 106754670
9 30077047 30187058 30297069
10 124176052 125276162 126376272
11 375067473 387058683 399049893
12 106700 107800 108900
13=1 77077 88088 99099
Table A.3: An IDR limiting 12-cycle in base-8 (octal) and base-9 (nonal) numerical system that has a correspondence in the base-10 (decimal) system.
An example of 41-cycle of IDR in the base-4 system
i 41-cycle elements (b=4)
1 13333032313332
2 103331022013323
3 1033301302213230
4 110113211113323
5 1100030330031000
8 1032123333130323
7 10323103333003230
8 1033010002210323
9 10323132002320230
10 1120211313121323
11 11012031110302200
12 10131123331221123
13 102310003330000230
14 10303310023321023
15 102322302031311330
16 3203111222022123
17 13021320103201212
18 100232210212113303
19 1010210022231011310
20 213102100030231203
21 1121300130032033121
22 3001203031002130332
23 10231023033103323
24 103221222131123130
25 11300030303000223
26 110100121212001200
27 101333303031000123
28 1023000100000333230
29 33010033330330023
30 1000123331322330
31 1231332112323
32 11130111110310
33 3223000001133
34 13200000011022
35 101211000011313
36 1020321000130020
37 220010332233213
38 1133003231303301
39 33311302233330
40 23312021321331
41 3333103233333
42=1 13333032313332
Table A.4: A base-4 (quaternary) limiting 41-cycle of IDR that does not have a correspondence in decimal system.